OpenStudy (aaronandyson):

eight metallic spheres each of radius 2 mm are melted and casted into a single sphere find the radius of the new metallic sphere @butterflydreamer

OpenStudy (butterflydreamer):

okayy so first step would be to find the total volume of the eight spheres.

OpenStudy (aaronandyson):

so i find V of one sphere and then times 8 it?

OpenStudy (aaronandyson):

V = volume

OpenStudy (butterflydreamer):

yyep!

OpenStudy (aaronandyson):

wait lemme calculate

OpenStudy (aaronandyson):

\[4/3 \times 22/7 \times 8? \]

OpenStudy (butterflydreamer):

mmm.. .you forgot the r^2. So V= 4/3* pi *r ^2

OpenStudy (aaronandyson):

hmm? i thought its r^3?

OpenStudy (butterflydreamer):

oh yeah LOL my bad! V = 4/3* pi * r^3 Then multiply that by 8

OpenStudy (aaronandyson):

4/3*22/7*64*8?

OpenStudy (aaronandyson):

2145.52 TIMES 8?

OpenStudy (butterflydreamer):

.................. no

OpenStudy (butterflydreamer):